For sample average = 1.5 and a μ = 1.52 and a s = 7 and a sample size n = 10, compute the t-score. Round your answers to 2 decimal places and be sure to include the negative (-) if necessary, in your answer. Enter your response in the space provided here.
Answer #1
sample mean 'x̄= 1.500
sample size   n= 10.00
sample std deviation s= 7.000
std error 'sx=s/√n= 2.2136
test stat t ='(x-μ)*√n/sx= -0.01